A 12.8-g piece of Styrofoam carries a net charge of -0.684 µC and is suspended in equilibrium above the center of a large, horizontal sheet of plastic that has a uniform charge density on its surface. What is the charge per unit area on the plastic sheet?
Answer #1

Solution:

Using the formula,

=> q * E = m*g at equilibrium

=> q * (σ/2eph) = m*g (eph stands for ephsilon not)

=> σ = 2 * eph * m * g / q

=> σ = 2 * 8.85 * 10^-12 * 12.8 * 10^-3 * 9.8 / (-0.684 * 10^-6)

=> σ = -3.246 * 10^-6

=> σ = 3.246 micro C/m^2